the locus of a point such that the two tangents drawn from it to parabola y^2=4ax are such that the slope of one os double the other... pleaseee help.!The final answer is y^2=9/2ax

Grade:11

1 Answers

Vaibhav
7 years ago
Let m1=m and m2=2m eqn of tangent of parabola=y=mx+a/m from y=4ax^2 a=a y=2m+a/2m eq1 y=mx+a/m eq2nd solving eq we get y=3mx put m=y/3x in eq 2 you get result